The incomes of junior executives in a large corporation are normally distributed with a standard deviation of $1,200. A cutback is pending, at which time those who earn less than $15,000 are to be discharged. If such a cut represents 10% of the junior executives, what is the current mean salary of the group of junior executives
Answer by Boreal(15235)

15000 is the 10th percentile of a distribution with unknown mean and sd of $1200
z(0.10)=(x-mean)/sd
-1.28=(15000-x)/1200
-1536=15000-x
x=$16,536 mean salary.